Barbara borrows $3000. She agrees to make monthly interest payments on the loan and will build up a sinking fund with monthly deposits to repay the principal with a single payment 19 months from now. If the interest being charged on the loan is j12 = 8.5% and the interest being earned on the sinking fund is j12 = 5.4%, what is the monthly cost of the debt for Barbara?
Answer #1
Loan amount 3000
Interst rate 8.50%
Monthly rate 0.7083%
Monthly interst payment 21.25
Future Amount of Sinking Fund 3000
Interst rate 5.40%
Monthly rate (i/12) 0.45%
Future value of Annuity is calculated as follows
FV=P*((1+r)^n-1/r)
FV= future value=3000
P=periodic payment
r= rate of interest
n= number of payments
3000=p*((1+0.0045)^19-1/0.0045)
3000=p*(1.089053-1/0.0045)
P=3000/19.78948
p=151.5957
Monthly cost =21.25+151.5957
$                                                                                  172.85